OpenStudy (anonymous):

Ignoring air resistance. One athlete has a vertical “jump” of 20 inches. What is the initial velocity required to jump this high?

OpenStudy (anonymous):

@AnimalAin can u help?

OpenStudy (anonymous):

Sure. We know that \[s(t)=-16t^2+v_0t+s_0\]works for projectile motion without aero drag. The ball player starts and ends on the floor, so s sub 0 = 0. That means\[s(t)=-16t^2+v_0t\]

OpenStudy (anonymous):

We know the t value at maximum height is \[t_{\max s}=\frac{-b}{2a}=\frac{v_0}{32}\]and that 20 inches is 5/3 feet, which is the maximum height.

OpenStudy (anonymous):

So, by substitution,\[\frac{5}{3}=-16(\frac{v_0}{32})^2+v_0(\frac{v_0}{32})\]Can you solve that for v0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hmm i'm trying to solve but im getting it wrong

OpenStudy (anonymous):

OK, lets keep going here.... If we simplify the RHS of my equation above, we will get\[\frac{5}{3}=\frac{v_0^2}{64} \implies v_0=\frac{8\sqrt{15}}{3} \approx 10.3~fps\]

OpenStudy (anonymous):

Hope that was helpful. Gotta go. Do math every day.

OpenStudy (anonymous):

ok thx